Manually calculate the compound interest on an investment of $12,500 at 7.5% interest, compounded semiannually, for 2 years. (Round value to the nearest cent.)

Answer by lwsshak3(11628)

Manually calculate the compound interest on an investment of $12,500 at 7.5% interest, compounded semiannually, for 2 years. (Round value to the nearest cent.)
***
compound interest formula: A=P(1+i)^n,P=initial investment, i=interest rate per compounding period, n=number of periods, A=amt after n-periods
..
For given problem:
P=12500
i=.075/2=0.0375
n=4
..
A=12500(1+.0375)^4
A=12500(1.0375)^4=14483.13
amt gained-initial investment=14483.13-12500.00=1983.13
Compound Interest=$1,983.13